Match the following aqueous solutions with the appropriate letter from the column on the right. 1. 0.13 m CoI2 A. Lowest freezing point 2. 0.10 m Cr(CH3COO)3 B. Second lowest freezing point 3. 0.16 m Mn(CH3COO)2 C. Third lowest freezing point 4. 0.44 m Glucose(nonelectrolyte) D. Highest freezing point
Answer #1

1. CoI2  --> Co2+ + 2 I-

for strong electrolyte van't hoff factor

i = number of ions

i = 1 + 2 or i = 3

Given molality = 0.13 m

Delta (Tf) = i *Kf *m

Put the value

Delta (Tf) = 3*Kf*0.13

Delta (Tf) = 0.39* Kf   ------(1)

2.

Cr(CH3COO)3 ---> Cr3+  + 3 CH3COO-

i = 1+3 or i = 4

molality = 0.10 m

Delta (Tf) = i *Kf *m

Delta (Tf) = 4*Kf *0.10

Delta (Tf) = 0.40*Kf ------(2)

3. Mn(CH3COO)2 ----> Mn2+ + 2 CH3COO-

i = 1+2 or i = 3

molality = 0.16 m

Delta (Tf) = i *Kf *m

Delta (Tf) = 3*Kf *0.16

Delta (Tf) = 0.48*Kf ----(3)

4.

Glucose will not dissolve into ions hence

i = 1

molality = 0.44 m

Delta (Tf) = i *Kf *m

Delta (Tf) = 1*Kf *0.44

Delta (Tf) = 0.44*Kf ---(4)

**** if delta( Tf ) will be high then freezing point will also be maximum( in negative).

From the above four equation

1. CoI2 A) lowest freezing point

2. Cr(CH3COO)3    B) second lowest freezing point

3. Mn(CH3COO)2   D) Highest freezing point

4. Glucose. C) third lowest freezing point